### v1

@@ Abstract
Predicting treatment non-response for anxiety and depression is challenging, in part because of sparse symptom assessments in real-world care. We examined whether passively captured, fine-grained emotions serve as linguistic markers of treatment outcomes by analyzing 12 weeks of de-identified teletherapy transcripts from 12,043 U.S. patients with moderate-to-severe anxiety and depression symptoms. A transformer-based small language model extracted patients’ emotions at the talk-turn level; a state-space model (VISTA) clustered subgroups based on emotion dynamics over time and produced temporal networks. Two groups emerged: an improving group (n=8,230) and a non-response group (n=3813) showing increased odds of symptom deterioration, and lower likelihood of clinically significant improvement. Temporal networks indicated that sadness and fear exerted most influence on emotion dynamics in non-responders, whereas improving patients showed balanced joy, sadness, and neutral expressions. Findings suggest that linguistic markers of emotional inflexibility can serve as scalable, interpretable, and theoretically grounded indicators for treatment risk stratification.
@@ 1 Introduction
Depression and anxiety represent a growing public health crisis with substantial emotional, social, and economic costs [1]. As the demand for care rises, digital health interventions offer an accessible alternative to traditional in-person treatment [2, 3]. Yet despite its promise, many participants fail to benefit, as with other mental health interventions. Predicting non-response is often hampered by a lack of consistent symptom assessments, and, in response to this challenge, there is increasing interest in predict treatment outcomes using behavioral or linguistic digital markers [4, 5]. These computational models could perform a critical role in identifying individuals at risk of non-response, and informing the allocation or reallocation of scarce therapeutic resources. However, for these methods to also be clinically useful, they must be grounded in a theoretical framework that highlights modifiable treatment targets.
@@
One promising avenue is the analysis of automatically-collected linguistic markers of emotional expression in psychotherapy. Prior studies have demonstrated that early use of positive sentiment, change talk, linguistic distancing, and markers of behavioral activation are associated with greater symptom improvements [6, 7, 8, 9, 10]. Other work has explored changes in emotional expression over the course of therapy, finding that an increase in positive sentiment and a decrease in first-person singular words throughout therapy were associated with symptom reductions [8, 11]. However, much of this work is limited by relying on lexicon- or rule-based approaches such as LIWC and VADER [6, 10, 8, 9], which cannot capture context-dependent meaning or linguistic nuance. Subsequent studies have applied deep learning models that can detect subtle emotional cues including recurrent neural networks (e.g., BiLSTM [7]), which capture sequential context, and transformer-based approaches (e.g. [11]), which capture long-range context. However, another key limitation across these studies is that they aggregate sentiment at the session or even multi-week level, a coarse granularity that misses clinically meaningful emotional fluctuations occurring within psychotherapy [4, 12, 13, 14].
@@
Parallel evidence from experience sampling and ecological momentary assessment (EMA) underscores the importance of short-timescale affect dynamics for treatment response. Some work suggests that baseline levels of positive and negative affect, or the ratio between positive and negative affect, are sufficient predictors of treatment response [15, 16]. However, many studies indicate that emotional fluctuations are important predictors of therapy outcomes. For example, [16] found that fluctuations in negative affect pre-treatment predicted early treatment response, whereas mean affect levels pre-treatment did not. Some work has explored the ways in which positive affect interacts with daily regulation processes, either by enhancing emotion regulation [17] or by suppressing subsequent negative affect [18]. In a complementary finding, [19] reported that lower emotional reactivity to both positive and negative daily events at baseline predicted greater depression symptom severity one month into treatment, suggesting that both the presence and functional significance of affective dynamics are important for treatment response.
@@
Part of the challenge lies in grounding investigations of emotion dynamics in theories that account for their complexity. Conventional perspectives on the role of emotion in psychotherapy often fall prey to the fallacy of uniform efficacy [20]; that is, they typically frame painful emotions, like fear or anger, as indicators of dysfunction, while positive states, such as joy and happiness, are viewed as uniform markers of mental health. Although the simplicity of this supposition is appealing, in their immediate moment-to-moment manifestations, emotions are essentially functional such that variations in both positive and negative emotions can be adaptive or maladaptive depending on the context in which they occur [21, 22, 23, 24]. A growing body of research has shown this to be the case as well for the strategic regulation of emotion [25, 26] (for review see [27]). Integrative theoretical models developed in response to this research, such as the regulatory flexibility framework [20, 28], provide a dynamic vehicle for understanding how individuals adaptively manage emotional experiences in the face of stress. This approach shifts attention from the simple presence of putatively healthy or unhealthy emotions to their skillful modulation in response to momentary shifts in situational demands. Regulatory flexibility thus highlights how balanced and varied emotional experience and expression may support adjustment, resilience, and therapeutic responsiveness, in addition to informing analytic methods that can identify more complex relationships between emotions and adjustment.
@@
Taken together, transcript-based sentiment analysis and EMA studies both highlight the relevance of emotional dynamics for therapy outcomes, but existing approaches either lack temporal precision (averaging sentiment across therapy sessions) or impose heavy participant burden (EMA). This gap points to the need for methods that capture fine-grained, in-session emotional fluctuations directly from message-based psychotherapy data. However most longitudinal clustering algorithms are unable to handle data that is both unequally spaced over time and imbalanced in number of observations [29]. As such, there is a need for longitudinal clustering algorithms that can handle irregularly sampled and varying time series to identify subgroups of patients that exhibit unique dynamic patters and links them to treatment outcomes.
@@
In the current study, we leverage a large psychotherapy cohort of over twelve thousand patients to identify emotion dynamics over treatment associated with outcomes. We first use a Small Language Model (SLM) [30] to extract utterance-level patients’ emotions from therapy transcripts. We then capture temporal flow across multiple emotions over treatment using VISTA (Varying and Irregular Sampling Time-series Analysis via State Space Models, see [31]), and test whether resulting clusters of emotional dynamics are prospectively associated with depression and anxiety outcomes. Finally, we construct cluster-specific temporal networks to interpret the underlying emotional dynamics and identify emotions with greater causal influence within each network, as shown in figure 1. By linking early patterns of emotional expressions to subsequent treatment response, this approach could support risk stratification and adaptive allocation of treatment resources.
@@ 2.1 Participants and Setting
The study sample comprised individuals in the United States seeking mental health treatment who enrolled in a digital mental teletherapy platform (www.talkspace.com) utilized by independent licensed clinicians. Access to this platform is available through employee assistance programs, direct self-enrollment, and coverage under individual insurance behavioral health benefits. The study consisted of individuals engaged in message-based psychotherapy [2] through the platform between 2016 and 2021. Prior to platform use, all patients and clinicians provided written consent to the use of their de-identified data for research purposes. Study procedures were approved by the NYU School of Medicine IRB (i23-00345).
@@ 2.1.1 Sample
Study eligibility required patients to have been diagnosed with depression or anxiety by their assigned licensed mental health provider and to have clinically significant symptoms at baseline, defined as scoring 10 or higher on the Patient Health Questionnaire-9 (PHQ-9) for depression and/or the Generalized Anxiety Disorder-7 (GAD-7) for anxiety [32, 33, 34]. Exclusion criteria consisted of comorbid psychosis, bipolar, and autism, as these conditions present unique clinical trajectories, or therapy transcripts deviating from symptoms assessment and typical engagement patterns.
@@ 2.2 Transcripts and Assessments
Patients were assigned to online “rooms” where they communicated with a licensed clinician. Patients could message their provider during office hours, typically 9 AM to 6 PM in the provider’s time zone, five days per week. For further details on message-based therapy, see Pullmann et al. [2]). We focused on the initial twelve weeks of treatment, consistent with psychotherapy standards for early treatment evaluation and response monitoring. Psychotherapy transcripts were generated and de-identified from the room logs by removing any personal identifiers, proper nouns, locations and dates.
@@ 2.3.1 Emotions Analysis
Psychotherapy transcripts were segmented into discrete patient talk turns, defined as all messages sent from a patient before a response from their provider. Each talk turn could be of any length and was analyzed separately. In order to detect emotions at the talk turn level, we leveraged the SLM DistilRoBERTa-base [35], which fine tunes the encoder-based transformer model RoBERTa [36] to identify emotions in English text data. This model has been applied in psycholinguistic analyses [37] and research examining the relationship between sentiment and mental health [38]. We used DistilRoBERTa-base to transform each talk turn into a vector of length seven, where each element ranges $[0,1]$ representing the strength in the text of a specific emotion (anger, disgust, fear, joy, neutral, sadness, and surprise).
@@ 2.3.2 Longitudinal Clusters of Emotions
We performed model-based longitudinal clustering of talk-turn emotion vectors over the 12-week treatment period using VISTA [31]. VISTA enables simultaneous parametric modeling and clustering of multivariate time series data with irregular and varying sampling rates, making it suitable for our scenario, as the number of talk-turns and their temporal frequency varied across patients. VISTA demonstrated superior performance compared to classical clustering methods and, when applied to psychological data, produced unsupervised clusters that corresponded with mental health information withheld from the model, supporting its use in situations without known “ground truth” labels for patient time series [31].
We provide a brief overview of VISTA. Given a set of $N$ unlabelled time series $Y^{i}=(y^{i}_{k})_{k=1,...,T_{i}}$ irregularly sampled at times $t_{1},\dots,t_{T_{i}}$ where $T_{i}$ denotes the number of samples for time series $i$, VISTA automatically assigns each of these to a certain group (or cluster), with the number $M$ of those groups being predefined. The determination of the clusters and the cluster assignment relies on a generative model whereby each time series in a specific cluster is assumed to arise as a realization of a linear stochastic dynamical system known as a linear Gaussian state space model [39] of the form:
@@
in which $l$ denotes the cluster to which the time series corresponds, $x_{k}$ is an underlying latent process generating the observations, and $A_{l,k}=\text{Id}+\Delta_{k}A_{l}$. Moreover, $\{w_{l,k}\}$, $\{v_{l,k}\}$, and $u_{l}$ are all Gaussian independent and identically distributed random variables with zero mean and respective covariance matrices $\Gamma_{l,k}=\Delta_{k}\Gamma_{l}$, $\Sigma_{l,k}=\frac{1}{\Delta_{k}}\Sigma_{l}$ and $P_{l}$, respectively, where $\Delta_{k}=t_{k}-t_{k-1}$. Therefore each cluster $l=1,\ldots,M$ is represented by the parameter set $\theta_{l}=\{\mu_{l},A_{l},C_{l},P_{l},\Sigma_{l},\Gamma_{l}\}$. The whole model behind VISTA can be then interpreted as an extension of the mixture of LGSSMs approach of [40]. More specifically, VISTA jointly and iteratively estimates both cluster assignments and cluster parameters based on the expectation-maximization (EM) algorithm. Upon convergence of the EM scheme, VISTA returns the cluster membership for each time series and the parameter set $\theta_{l}=\{\mu_{l},A_{l},C_{l},P_{l},\Sigma_{l},\Gamma_{l}\}$ for each cluster $l=1,\dots,M$. These parameters provide interpretability to the estimated clusters as they can be leveraged to simulate a “typical” trajectory for each group by taking (1) without noise terms removed. For a more in depth description of the approach, see Brindle et al. [31].
We employed VISTA to identify heterogeneous clusters based on the trajectories of seven emotions obtained from patient talk-turn over twelve weeks of treatment. VISTA analyzed emotion vectors and their associated timestamps to produce unsupervised cluster assignments, along with a parameterized model for each cluster that describes the temporal evolution of sentiment. We examined baseline symptom-level differences in anxiety and depression symptoms between clusters using the Mann-Whitney U test [41] and adjusted p-values for multiple testing using Bonferroni correction [42]. Employing logistic regression, we then associated patient-level cluster membership with clinical outcomes over 12 weeks of treatment. We defined patient outcomes of clinically significant change [43] in either PHQ-9 or GAD-7 as moving below the clinical threshold (score of <10<10) and improving at least 5 points and deterioration as the worsening of symptoms by 5 or more points.
@@ 2.3.3 Temporal Networks of Emotions
After defining clusters of temporal dynamics among the seven emotions and their association with clinical improvement and non-response, we identified the most central emotions driving these dynamics based on psychological network theory [44]. We estimated temporal networks from patients’ talk-turn emotions dynamics as a second step of the VISTA clustering procedure. Notably, characterizing how observed variables influence one another differently across clusters and throughout time evolution is a more challenging task than simply defining their longitudinal course. From (1), we observe that the system does not directly relate $y_{k}$ and $y_{k-1}$, but rather models interactions at the level of the latent variables via the matrices $A_{l,k}$. In order to approximate the influence of $y_{k-1}$ on $y_{k}$, we observe that $x_{k}\approx A_{l,k}x_{k-1}$ and $y_{k}\approx C_{l}x_{k}$, yielding $y_{k}\approx C_{l}A_{l,k}C^{+}_{l}y_{k-1}$. $C^{+}_{l}$ is the Moore-Penrose pseudoinverse [45], used because $C_{l}$ is only a square matrix when $x$ and $y$ are of equal dimension, while in practice, a latent variable with higher dimension than the observed typically gives more meaningful results. This approximation allows us to consider the evolution of the observed variable with a simple linear equation for each cluster, shown in (2):
@@
For analysis at the cluster level, rather than at the individual level, we fix $\Delta_{k}$, which leads to a linear update equation relating $y_{k-1}$ to $y_{k}$ independent of $k$. In our setting, we set $\Delta_{k}$ to represent one week due to the assessment period being broken into three week intervals and the assessments themselves phrasing questions in terms of weeks. As we are interested in learning how each individual emotion in the talk turn influences the time evolution of the others throughout the clusters of patients returned by VISTA, we construct a temporal network from the matrix $C_{l}(\text{Id}+\Delta_{k}A_{l})C_{l}^{+}$. From the temporal network, we gain insight into which specific emotions dynamic expressed in patient transcripts are associated with clinical outcomes by examining the centrality of each emotion [44]. Specifically, we examine the Expected Influence [46] in the longitudinal network and compare centrality rankings [47] of the most central emotions in each temporal network.

### v2

@@ Abstract
Predicting treatment non-response for anxiety and depression is challenging, in part because of sparse symptom assessments in real-world care. We examined whether passively captured, fine-grained emotions serve as linguistic markers of treatment outcomes by analyzing 12 weeks of de-identified teletherapy transcripts from 12,043 U.S. patients with moderate-to-severe anxiety and depression symptoms. A transformer-based small language model extracted patients’ emotions at the talk-turn level; a state-space model (VISTA-SSM) clustered subgroups based on emotion dynamics over time and produced temporal networks. Two groups emerged: an improving group (n=8,230) and a non-response group (n=3,813) showing increased odds of symptom deterioration, and lower likelihood of clinically significant improvement. Temporal networks indicated that sadness and fear exerted most influence on emotion dynamics in non-responders, whereas improving patients showed balanced joy, sadness, and neutral expressions. Findings suggest that linguistic markers of emotional inflexibility can serve as scalable, interpretable, and theoretically grounded indicators for treatment risk stratification.
@@ 1 Introduction
Depression and anxiety represent a growing public health crisis with substantial emotional, social, and economic costs [counts2023psychological]. As the demand for care rises, digital health interventions offer an accessible alternative to traditional in-person treatment [pullmann2025message, hull2020two]. Yet despite its promise, many participants fail to benefit, as with other mental health interventions. Predicting non-response is often hampered by a lack of consistent symptom assessments, and, in response to this challenge, there is increasing interest in predicting treatment outcomes using behavioral digital markers [doherty2025passive]. Computational models could perform a critical role in identifying individuals at risk of non-response, and informing the allocation or reallocation of scarce therapeutic resources. However, for these methods to also be clinically useful, they must be grounded in a theoretical framework that highlights interpretable treatment targets.
@@
One promising avenue is the analysis of automatically-collected linguistic markers of emotional expression [malgaroli2023natural]. Prior studies have demonstrated that early use of positive sentiment, change talk, linguistic distancing, and markers of behavioral activation are associated with greater symptom improvements [Burkhardt2021, Ewbank2021, Shapira2022, Malgaroli2024, Nook2022]. Other work has explored changes in emotional expression over the course of treatment, finding that an increase in positive sentiment and a decrease in first-person singular words throughout therapy were associated with symptom reductions [Shapira2022, Eberhardt2025]. However, much of this work is limited by relying on lexicon- or rule-based approaches such as LIWC and VADER [Burkhardt2021, Nook2022, Shapira2022, Malgaroli2024], which cannot capture context-dependent meaning or linguistic nuance. Subsequent studies have applied deep learning models that can detect subtle emotional cues including recurrent neural networks that capture sequential dependencies (e.g., BiLSTM [Ewbank2021]), and transformer-based approaches that capture long-range context (e.g. [Eberhardt2025]). However, another key limitation across these studies is that they aggregate sentiment at the week level, a coarse temporal granularity that misses clinically meaningful emotional fluctuations occurring within treatment [malgaroli2023natural, Luo2025, Paz2024].
@@
Parallel evidence from experience sampling and ecological momentary assessment (EMA) underscores the importance of short-timescale affect dynamics for treatment response. While some work suggests that baseline levels of positive and negative affect, or the ratio between them, are sufficient predictors of treatment response [Forbes2012, Husen2016], many studies indicate that emotional fluctuations carry additional predictive value. For example, Husen and colleagues [Husen2016] found that fluctuations in negative affect pre-treatment predicted early treatment response, whereas mean affect levels did not. Other work has explored how positive affect interacts with daily regulatory processes, either by enhancing emotion regulation [Hehlmann2024] or by suppressing subsequent negative affect [Wichers2012]. In a complementary finding, Peeters and colleagues [Peeters2010] reported that lower emotional reactivity to both positive and negative daily events at baseline predicted greater depression symptom severity one month into treatment, suggesting that both the presence and functional significance of affective dynamics are important for treatment response.
@@
Part of the challenge lies in grounding investigations of emotion dynamics in theories that account for their complexity. Conventional perspectives on the role of emotion in treatment often fall prey to the fallacy of uniform efficacy [bonanno2013regulatory]; that is, they typically frame painful emotions, like fear or anger, as indicators of dysfunction, while positive states, such as joy and happiness, are viewed as uniform markers of mental health. Although the simplicity of this supposition is appealing, in their immediate moment-to-moment manifestations, emotions are essentially functional such that variations in both positive and negative emotions can be adaptive or maladaptive depending on the context in which they occur [coifman2016context, coifman2010distress, gruber2011dark]. A growing body of research has shown this to be the case as well for the strategic regulation of emotion [bonanno2004importance, troy2013person] (for review see [bonanno2023resilience]). Integrative theoretical models developed in response to this research, such as the regulatory flexibility framework [bonanno2013regulatory, bonanno2021resilience], provide a dynamic vehicle for understanding how individuals adaptively manage emotional experiences in the face of stress. This approach shifts attention from the simple presence of putatively healthy or unhealthy emotions to their skillful modulation in response to momentary shifts in situational demands. Regulatory flexibility thus highlights how balanced and varied emotional experience and expression may support adjustment, resilience, and therapeutic responsiveness. Critically, it motivates analytic methods that can capture dynamic relationships between emotions and adjustment.
@@
Taken together, transcript-based sentiment analysis and EMA studies both underscore the relevance of emotional dynamics for treatment outcomes, yet each approach has limitations: linguistic analyses typically aggregate across sessions, while EMA imposes heavy participant burden. This gap points to the need for methods that capture fine-grained emotional fluctuations directly. A further methodological challenge is that most longitudinal clustering algorithms are unable to handle data that is both unequally spaced over time and imbalanced in number of observations [lu2024clustering], necessitating specialized approaches to identify patient subgroups exhibiting distinct dynamic patterns and linking them to treatment outcomes.
@@
In the current study, we leverage a large psychotherapy cohort of over twelve thousand patients to identify emotion dynamics over treatment associated with outcomes. We first use a Small Language Model (SLM) [belcak2025small] to extract utterance-level patients’ emotions from therapy transcripts. We then capture temporal flow across multiple emotions over treatment using VISTA-SSM (Varying and Irregular Sampling Time-series Analysis via State Space Models, see [brindle2024vista]), and test whether resulting clusters of emotional dynamics are prospectively associated with depression and anxiety outcomes. Finally, we construct cluster-specific temporal networks to interpret the underlying emotional dynamics and identify emotions with greater causal influence within each network, as shown in figure 1. By linking early patterns of emotional expressions to subsequent treatment response, this approach could support risk stratification and adaptive allocation of treatment resources.
@@ 2.1 Participants and Setting
The study sample comprised individuals in the United States seeking mental health treatment who enrolled in a digital mental teletherapy platform (www.talkspace.com) utilized by independent licensed clinicians. Access to this platform is available through employee assistance programs, direct self-enrollment, and coverage under individual insurance behavioral health benefits. The study consisted of individuals engaged in message-based psychotherapy [pullmann2025message] through the platform between 2016 and 2021. Prior to platform use, all patients and clinicians provided written consent to the use of their de-identified data for research purposes. Study procedures were approved by the NYU School of Medicine IRB (i23-00345).
@@ 2.1.1 Sample
Study eligibility required patients to have been diagnosed with depression or anxiety by their assigned licensed mental health provider and to have clinically significant symptoms at baseline, defined as scoring 10 or higher on the Patient Health Questionnaire-9 (PHQ-9) for depression and/or the Generalized Anxiety Disorder-7 (GAD-7) for anxiety [kroenke2001phq, kroenke2007anxiety]. Exclusion criteria consisted of comorbid psychosis, bipolar, and autism, as these conditions present unique clinical trajectories, or therapy transcripts deviating from symptoms assessment and typical engagement patterns.
@@ 2.2 Transcripts and Assessments
Patients were assigned to online “rooms” where they communicated with a licensed clinician. Patients could message their provider during office hours, typically 9 AM to 6 PM in the provider’s time zone, five days per week. For further details on message-based therapy, see Pullmann et al. [pullmann2025message]). We focused on the initial twelve weeks of treatment, consistent with psychotherapy standards for early treatment evaluation and response monitoring. Psychotherapy transcripts were generated and de-identified from the room logs by removing any personal identifiers, proper nouns, locations and dates.
@@ 2.3.1 Emotions Analysis
Psychotherapy transcripts were segmented into discrete patient talk turns, defined as all messages sent from a patient before a response from their provider. Each talk turn could be of any length and was analyzed separately. In order to detect emotions at the talk turn level, we leveraged the SLM DistilRoBERTa-base [hartmann2022emotionenglish], which fine tunes the encoder-based transformer model RoBERTa [liu2019roberta] to identify emotions in English text data. This model has been applied in psycholinguistic analyses [butt2022goes] and research examining the relationship between sentiment and mental health [pathak2025sentiment]. We used DistilRoBERTa-base to transform each talk turn into a vector of length seven, where each element ranges $[0,1]$ representing the strength in the text of a specific emotion (anger, disgust, fear, joy, neutral, sadness, and surprise).
@@ 2.3.2 Longitudinal Clusters of Emotions
We performed model-based longitudinal clustering of talk-turn emotion vectors over the 12-week treatment period using VISTA [brindle2024vista]. VISTA enables simultaneous parametric modeling and clustering of multivariate time series data with irregular and varying sampling rates, making it suitable for our scenario, as the number of talk-turns and their temporal frequency varied across patients. VISTA demonstrated superior performance compared to classical clustering methods and, when applied to psychological data, produced unsupervised clusters that corresponded with mental health information withheld from the model, supporting its use in situations without known “ground truth” labels for patient time series [brindle2024vista].
We provide a brief overview of VISTA. Given a set of $N$ unlabelled time series $Y^{i}=(y^{i}_{k})_{k=1,...,T_{i}}$ irregularly sampled at times $t_{1},\dots,t_{T_{i}}$ where $T_{i}$ denotes the number of samples for time series $i$, VISTA automatically assigns each of these to a certain group (or cluster), with the number $M$ of those groups being predefined. The determination of the clusters and the cluster assignment relies on a generative model whereby each time series in a specific cluster is assumed to arise as a realization of a linear stochastic dynamical system known as a linear Gaussian state space model [ho1964bayesian] of the form:
@@
in which $l$ denotes the cluster to which the time series corresponds, $x_{k}$ is an underlying latent process generating the observations, and $A_{l,k}=\text{Id}+\Delta_{k}A_{l}$. Moreover, $\{w_{l,k}\}$, $\{v_{l,k}\}$, and $u_{l}$ are all Gaussian independent and identically distributed random variables with zero mean and respective covariance matrices $\Gamma_{l,k}=\Delta_{k}\Gamma_{l}$, $\Sigma_{l,k}=\frac{1}{\Delta_{k}}\Sigma_{l}$ and $P_{l}$, respectively, where $\Delta_{k}=t_{k}-t_{k-1}$. Therefore each cluster $l=1,\ldots,M$ is represented by the parameter set $\theta_{l}=\{\mu_{l},A_{l},C_{l},P_{l},\Sigma_{l},\Gamma_{l}\}$. The whole model behind VISTA can be then interpreted as an extension of the mixture of LGSSMs approach of [umatani2023time]. More specifically, VISTA jointly and iteratively estimates both cluster assignments and cluster parameters based on the expectation-maximization (EM) algorithm. Upon convergence of the EM scheme, VISTA returns the cluster membership for each time series and the parameter set $\theta_{l}=\{\mu_{l},A_{l},C_{l},P_{l},\Sigma_{l},\Gamma_{l}\}$ for each cluster $l=1,\dots,M$. These parameters provide interpretability to the estimated clusters as they can be leveraged to simulate a “typical” trajectory for each group by taking (1) without noise terms removed. For a more in depth description of the approach, see Brindle et al. [brindle2024vista].
We employed VISTA to identify heterogeneous clusters based on the trajectories of seven emotions obtained from patient talk-turn over twelve weeks of treatment. VISTA analyzed emotion vectors and their associated timestamps to produce unsupervised cluster assignments, along with a parameterized model for each cluster that describes the temporal evolution of sentiment. We examined baseline symptom-level differences in anxiety and depression symptoms between clusters using the Mann-Whitney U test [mann1947test] and adjusted p-values for multiple testing using Bonferroni correction [bonferroni1936teoria]. Employing logistic regression, we then associated patient-level cluster membership with clinical outcomes over 12 weeks of treatment. We defined patient outcomes of clinically significant change [jacobson1992clinical] in either PHQ-9 or GAD-7 as moving below the clinical threshold (score of <10<10) and improving at least 5 points and deterioration as the worsening of symptoms by 5 or more points.
@@ 2.3.3 Temporal Networks of Emotions
After defining clusters of temporal dynamics among the seven emotions and their association with clinical improvement and non-response, we identified the most central emotions driving these dynamics based on psychological network theory [borsboom2021network]. We estimated temporal networks from patients’ talk-turn emotions dynamics as a second step of the VISTA clustering procedure. Notably, characterizing how observed variables influence one another differently across clusters and throughout time evolution is a more challenging task than simply defining their longitudinal course. From (1), we observe that the system does not directly relate $y_{k}$ and $y_{k-1}$, but rather models interactions at the level of the latent variables via the matrices $A_{l,k}$. In order to approximate the influence of $y_{k-1}$ on $y_{k}$, we observe that $x_{k}\approx A_{l,k}x_{k-1}$ and $y_{k}\approx C_{l}x_{k}$, yielding $y_{k}\approx C_{l}A_{l,k}C^{+}_{l}y_{k-1}$. $C^{+}_{l}$ is the Moore-Penrose pseudoinverse [ben2003generalized], used because $C_{l}$ is only a square matrix when $x$ and $y$ are of equal dimension, while in practice, a latent variable with higher dimension than the observed typically gives more meaningful results. This approximation allows us to consider the evolution of the observed variable with a simple linear equation for each cluster, shown in (2):
@@
For analysis at the cluster level, rather than at the individual level, we fix $\Delta_{k}$, which leads to a linear update equation relating $y_{k-1}$ to $y_{k}$ independent of $k$. In our setting, we set $\Delta_{k}$ to represent one week due to the assessment period being broken into three week intervals and the assessments themselves phrasing questions in terms of weeks. As we are interested in learning how each individual emotion in the talk turn influences the time evolution of the others throughout the clusters of patients returned by VISTA, we construct a temporal network from the matrix $C_{l}(\text{Id}+\Delta_{k}A_{l})C_{l}^{+}$. From the temporal network, we gain insight into which specific emotions dynamic expressed in patient transcripts are associated with clinical outcomes by examining the centrality of each emotion [borsboom2021network]. Specifically, we examine the Expected Influence [robinaugh2016identifying] in the longitudinal network and compare centrality rankings [malgaroli2021networks] of the most central emotions in each temporal network.